# A Survey of Reinforcement Learning For Economics

### 09 Mar 2026 v1

@@ A Survey of Reinforcement Learning For Economics
This survey (re)introduces reinforcement learning methods to economists. The curse of dimensionality limits how far exact dynamic programming can be effectively applied, forcing us to rely on suitably “small” problems or our ability to convert “big” problems into smaller ones. While this reduction has been sufficient for many classical applications, a growing class of economic models resists such reduction. Reinforcement learning algorithms offer a natural, sample-based extension of dynamic programming, extending tractability to problems with high-dimensional states, continuous actions, and strategic interactions. I review the theory connecting classical planning to modern learning algorithms and demonstrate their mechanics through simulated examples in pricing, inventory control, strategic games, and preference elicitation. I also examine the practical vulnerabilities of these algorithms, noting their brittleness, sample inefficiency, sensitivity to hyperparameters, and the absence of global convergence guarantees outside of tabular settings. The successes of reinforcement learning remain strictly bounded by these constraints, as well as a reliance on accurate simulators. When guided by economic structure, reinforcement learning provides a remarkably flexible framework. It stands as an imperfect, but promising, addition to the computational economist’s toolkit. A companion survey (Rust and Rawat, 2026b) covers the inverse problem of inferring preferences from observed behavior.
@@ 1 Introduction
Reduction of average Bellman errors gives dynamic programming a geometric rate of convergence to the optimal solution, while the incremental updates and reduction of sampled Bellman errors, when combined with “sufficient exploration” of the state-action space, gives reinforcement learning only sublinear convergence guarantees. This is however, quite sufficient in practise, the scalability attained by only sampling transitions and making incremental updates more than makes up for the slower rates of convergence (and brittleness).
These approximation methods sacrifice theoretical guarantees. RL algorithms lack the convergence assurances of exact dynamic programming. They exhibit sensitivity to hyperparameters and initialization. They can converge to suboptimal policies without diagnostic indication. This survey presents reinforcement learning as a computationally flexible framework while acknowledging its methodological limitations.
@@
This survey adopts a unified notation for concepts that developed independently across these traditions. A Markov decision process consists of a state space $\mathcal{S}$, an action space $\mathcal{A}$, a transition kernel $P(s^{\prime}|s,a)$ giving the probability of reaching state $s^{\prime}$ from state $s$ under action $a$, a reward function $r(s,a)$, and a discount factor $\gamma\in[0,1)$. A policy $\pi(a|s)$ specifies the probability of selecting action $a$ in state $s$; a deterministic policy is written $\pi(s)$. The state-value function $V^{\pi}(s)=\mathbb{E}_{\pi}[\sum_{t=0}^{\infty}\gamma^{t}r(s_{t},a_{t})\mid s_{0}=s]$ measures expected discounted return from state $s$ under policy $\pi$; the action-value function $Q^{\pi}(s,a)$ conditions additionally on the first action. The optimal value functions $V^{*}$ and $Q^{*}$ correspond to the policy that maximizes expected return. Value functions are often approximated by parameterized functions: $V(s;\mathbf{w})$ for linear approximations with weight vector $\mathbf{w}$, and $V(s;\theta)$ for general parameterizations including neural networks. Updates proceed by a learning rate $\alpha>0$. The temporal difference error $\delta_{t}=r_{t+1}+\gamma V(s_{t+1})-V(s_{t})$ measures the discrepancy between successive value estimates and drives most learning algorithms in this survey.
@@ 2.2 Board Games
The linear parameterization compresses the value function from $10^{20}$ table entries to dozens of weights, the essential response to the curse of dimensionality. Samuel’s architecture — minimax search to depth $H$ with the learned evaluation function scoring leaves — is an early instance of rollout, where tree search simulates forward, truncating at $H$ and substituting $V(s)$ for exact continuation values. [^5] Bertsekas (2021) interprets this as approximate dynamic programming, where offline training (learning $V$) combined with online planning (tree search) implements a Newton-like step for the Bellman equation. The conceptual apparatus of modern game-playing AI — self-play, evaluation learning, tree search, and function approximation — was present in the 1950s.
@@ 2.3 Optimal Control
Blackwell (1965) established the measure-theoretic foundations for discounted dynamic programming with general state and action spaces. He proved that the Bellman operator $T$ defined by $Tu(s)=\sup_{a}[r(s,a)+\gamma\int u(s^{\prime})P(ds^{\prime}|s,a)]$ is a contraction with modulus $\gamma$: $\|Tu-Tv\|_{\infty}\leq\gamma\|u-v\|_{\infty}$. Banach’s fixed-point theorem then guarantees a unique bounded solution $V^{*}$ to the Bellman equation, with $\|V_{k}-V^{*}\|_{\infty}\leq\gamma^{k}\|V_{0}-V^{*}\|_{\infty}$ under value iteration. The central result concerns stationary policies, which use the same decision rule $f:\mathcal{S}\to\mathcal{A}$ at every period regardless of history. Blackwell proved that if the action space is finite, there exists an optimal stationary policy. For countable action spaces, $\epsilon$-optimal stationary policies exist for every $\epsilon>0$. These results justify the focus on memoryless policies, since optimal behavior depends only on the current state, not on the history of past states and actions. The Bellman equation, Howard’s policy iteration, and Blackwell’s existence theorems constitute the planning framework. Given complete knowledge of $P$ and $r$, these methods compute optimal policies exactly. The challenge of learning without such knowledge is the central problem of reinforcement learning.
@@ 3 Reinforcement Learning Algorithms
In this section I introduce important reinforcement learning algorithms while pushing the theoretical results to the next chapter. The central problem is how do we move from estiamating $V_{\pi}$ and $V_{*}$ when we do not have a model.
@@ 3.1.1 Monte Carlo Estimation
When $P(s^{\prime}|s,a)$ and $r(s,a)$ are unknown, the obvious approach is to use Monte Carlo (MC) to approximate them. These methods estimate value functions from sampled episodes $(s_{0},a_{0},r_{1},s_{1},a_{1},r_{2},\ldots,s_{T})$. The realized return from state $s_{t}$ is:
@@ 3.1.2 Sutton (1988)
The target $r_{t+1}+\gamma V(s_{t+1})$ depends on one random reward and one random transition, so its variance is low. The cost is bias: $V(s_{t+1})$ is the agent’s current estimate, not the true value. This is “bootstrapping”. [^8] Bootstrapping means updating value estimates using other value estimates rather than observed returns. As $V$ improves, the bias shrinks; the low variance persists regardless. Sutton demonstrated this tradeoff on a five-state random walk where TD(0) converged faster than Monte Carlo with less data.
The general TD($\lambda$) update interpolates between these extremes through an eligibility trace: [^9] The eligibility trace records which states were recently visited, allowing credit assignment to propagate backward in time. States visited more recently receive stronger updates when the TD error is observed.
@@ 3.1.3 Watkins (1989)
Q-learning achieves this via incremental updates:
@@
Q-learning converges to $Q^{*}$ under standard regularity conditions (Section 4.2). The maximization over $a^{\prime}$ makes Q-learning off-policy: [^11] Off-policy learning means the agent learns about a policy different from the one generating data. Q-learning learns about the greedy policy $\pi^{*}(s)=\mathop{\it argmax}_{a}Q(s,a)$ while collecting data with an exploratory policy. This exploratory policy needs only to be sufficiently “exploratory” and admits a wide range of policies; including a fully random policy. the update target uses the greedy action at the next state regardless of the action actually taken. Therefore the agent can follow an $\varepsilon$-greedy exploration strategy or even a fully random policy, while learning about the optimal policy directly.
@@ 3.1.4 Williams (1992)
Value-based methods learn an action-value function and derive a policy from it. Williams (1992) derived an alternative, policy gradients, that optimize the policy directly by gradient ascent on expected return:
@@
Consider a robot arm that must apply a continuous torque $a\in\mathbb{R}$ to reach a target angle. Q-learning requires computing $\max_{a}Q(s,a)$ at every update, which becomes a nested optimization problem  [^12] In continuous action spaces, $\max_{a}Q(s,a)$ has no closed-form solution in general and must be solved numerically at every Bellman update. Discretizing a $d$-dimensional action space on a grid of $m$ points per dimension costs $O(m^{d})$ evaluations per update step. when the action space is continuous. A policy gradient method sidesteps the issue. Parameterize the policy as a Gaussian $\pi_{\theta}(a|s)=\mathcal{N}(\mu_{\theta}(s),\,\sigma_{\theta}^{2}(s))$, [^13] The Gaussian distribution $\mathcal{N}(\mu,\sigma^{2})$ has density $(2\pi\sigma^{2})^{-1/2}\exp(-(a-\mu)^{2}/2\sigma^{2})$. Here $\mu_{\theta}(s)$ and $\sigma_{\theta}(s)$ are neural network outputs parameterizing the policy mean and standard deviation. sample an action, observe the return, and update $\theta$ by REINFORCE. No maximization over actions is required. Virtually all continuous-control results in reinforcement learning descend from the policy gradient framework rather than from Q-learning for this reason.

### 15 Mar 2026 v2

@@ A Survey of Reinforcement Learning For Economics
This survey (re)introduces reinforcement learning methods to economists. The curse of dimensionality limits how far exact dynamic programming can be effectively applied, forcing us to rely on suitably “small” problems or our ability to convert “big” problems into smaller ones. While this reduction has been sufficient for many classical applications, a growing class of economic models resists such reduction. Reinforcement learning algorithms offer a natural, sample-based extension of dynamic programming, extending tractability to problems with high-dimensional states, continuous actions, and strategic interactions. I review the theory connecting classical planning to modern learning algorithms and demonstrate their mechanics through simulated examples in pricing, inventory control, strategic games, and preference elicitation. I also examine the practical vulnerabilities of these algorithms, noting their brittleness, sample inefficiency, sensitivity to hyperparameters, and the absence of global convergence guarantees outside of tabular settings. The successes of reinforcement learning remain strictly bounded by these constraints, as well as a reliance on accurate simulators. That said, when guided by economic structure, reinforcement learning provides a flexible and innovative framework. It stands as an imperfect, but promising, addition to the computational economist’s toolkit. A companion survey (Rust and Rawat, 2026b) covers the inverse problem of inferring preferences from observed behavior. All simulation code is publicly available. [^1] https://github.com/rawatpranjal/survey-of-reinforcement-learning-in-economics
@@ 1 Introduction
Reduction of average Bellman errors gives dynamic programming a geometric rate of convergence to the optimal solution, while the incremental updates and reduction of sampled Bellman errors, when combined with “sufficient exploration” of the state-action space, gives reinforcement learning only sublinear convergence guarantees. This is however, quite sufficient in practise, the scalability attained by only sampling transitions and making incremental updates more than makes up for the slower rates of convergence (and brittleness). These approximation methods sacrifice theoretical guarantees. RL algorithms lack the convergence assurances of exact dynamic programming. They exhibit sensitivity to hyperparameters and initialization. They can converge to suboptimal policies without diagnostic indication. This survey presents reinforcement learning as a computationally flexible framework while acknowledging its methodological limitations.
Theory in reinforcement learning trails empirical success, often by years; convergence guarantees, sample complexity bounds, and approximation error characterizations typically arrive after practitioners have demonstrated that an algorithm works. The theoretical insights that eventually follow tend to be deep and structural, and the empirical frontier is itself a productive research frontier. Experiments are brittle, conducted on benchmark environments that are stylized approximations of deployment settings. These benchmarks nonetheless serve a critical coordination function, aligning research effort, enabling reproducible comparison, and exposing failure modes that motivate new theory. Details matter disproportionately in reinforcement learning; small implementation choices can determine whether an algorithm converges or diverges, and the practice of releasing code and documenting hyperparameters, seeds, and preprocessing has proven essential to progress.
@@
2 The Language of Reinforcement Learning
RL organizes its core problems along axes that do not map directly onto econometric categories. Prediction refers to estimating $V^{\pi}(s)$ or $Q^{\pi}(s,a)$ for a fixed policy $\pi$ (policy evaluation), not forecasting observable variables. Control refers to finding the policy $\pi^{*}$ that maximizes expected discounted return (policy optimization), not the inclusion of regressors. Because prediction concerns evaluating a specific policy, a closely related question is whether the data was generated by the policy being studied. On-policy methods evaluate and improve the same policy that generates the data. Off-policy methods learn about a target policy $\pi$ from data generated by a different behavioral policy $\mu$. This distinction is central to causal inference (Section 13), where off-policy evaluation is precisely counterfactual policy evaluation.
A separate question concerns how the algorithm accesses experience. Online RL learns while interacting with the environment, collecting new data as a consequence of its own actions. Offline RL (also called batch RL) learns exclusively from a fixed dataset of previously collected transitions, with no ability to gather additional samples. The offline setting is closer to standard empirical work, where the dataset is given and the analyst cannot run new experiments; the online setting corresponds more closely to adaptive experimental design or sequential decision problems. Note that “online” in RL carries no timing constraint; it means only that the agent generates fresh experience. Real-time RL, by contrast, imposes hard deadlines on the perception-action loop, as in robotics or mechatronics. Every real-time RL system is online, but most online RL (games, recommender systems, RLHF) is not real-time. Orthogonal to data access is what the algorithm does with that data. A model-based RL algorithm explicitly learns or is given the transition kernel $P(s^{\prime}|s,a)$ and reward function $r(s,a)$, then plans using these learned dynamics. A model-free algorithm learns value functions or policies directly from sampled transitions without ever constructing an explicit dynamics model. In economics, “model-based” connotes structural estimation, the specification of a complete economic model with behavioral assumptions, equilibrium conditions, and identified parameters. An RL algorithm can be “model-based” in the RL sense (it learns $P$ and $r$) while being entirely “reduced-form” in the economics sense (it makes no behavioral or equilibrium assumptions). Conversely, a structural econometric model solved by simulation methods is “model-based” in the economics sense but may use “model-free” RL algorithms internally to approximate the value function.
Bandits sit naturally in this taxonomy. A multi-armed or contextual bandit is a one-step MDP: the agent observes context (state), selects an arm (action), receives an immediate reward, and the episode ends; actions do not affect future states. Bandits are therefore a special case of online RL with horizon one and no transition dynamics. They are not inherently real-time; a bandit algorithm can run under tight latency constraints, but that is an engineering requirement, not a defining property. In practice, bandits are the tool of choice when the next context is approximately independent of the current action and the problem reduces to exploration-exploitation without long-horizon credit assignment, as in dynamic pricing, ad placement, and clinical trials. Offline variants also exist: logged-bandit feedback methods learn from historically collected action-reward pairs without further interaction.
Overlapping Terminology
Perhaps the most consequential divergence is the word model. In structural econometrics, a model is a complete specification of preferences, technology, beliefs, and an equilibrium concept; estimating it means recovering the primitives that rationalize observed behavior. In RL, “model” refers exclusively to the pair $(P,r)$, the transition dynamics and reward function of the MDP, with no implications about objectives, strategic interaction, or equilibrium. The environment absorbs some of this gap: in RL it is a formal object encompassing everything outside the agent (the DGP, other agents, market clearing conditions), whereas in economics “environment” refers more loosely to market structure or institutional rules. [^2] A third source of confusion is generative model. In machine learning broadly, this means a model of the joint distribution $P(X,Y)$ or a synthetic data generator (GANs, diffusion models). In RL theory, a generative model is a simulator oracle that, given any $(s,a)$, returns a sample $s^{\prime}\sim P(\cdot|s,a)$ and reward $r(s,a)$; the usage implies random-access simulation, stronger than sequential online interaction but weaker than knowing $P$ analytically.
The terms for what is observed and measured diverge just as sharply. A return in finance is the yield on an investment, typically expressed as a percentage gain; in RL, the return $G_{t}=\sum_{k=0}^{\infty}\gamma^{k}r_{t+k+1}$ is the discounted sum of future rewards from time $t$ onward, the random variable whose expectation defines the value function. The RL usage is closer to what economists call the “present discounted value” of a stream of payoffs. A single complete sequence of interactions from an initial state to termination, $(s_{0},a_{0},r_{1},s_{1},\ldots,s_{T})$, is called an episode (synonymously, trajectory or rollout). [^3] The closest econometric analogs are a single panel unit’s time series, a realization of a stochastic process, or one “history” in a dynamic model. Where an econometrician speaks of the outcome, meaning the dependent variable $Y$ in a regression, RL has no single analog: the reward $r_{t}$ is the per-period outcome, the return $G_{t}$ is the cumulative outcome, and the value function $V^{\pi}(s)$ is the expected cumulative outcome conditional on state.
Several methodological terms carry misleading surface-level overlap. Bootstrapping in econometrics refers to Efron’s resampling method (Efron, 1979); in RL, it means updating a value estimate using another value estimate rather than a complete realized return, as when a TD algorithm uses the target $r_{t+1}+\gamma V(s_{t+1})$ that depends on the current, uncertain estimate $V(s_{t+1})$ (Sutton, 1988). Learning in economic theory (Bayesian learning, adaptive expectations) refers to agents updating beliefs about unknown parameters of their environment; in RL, learning refers to the algorithmic process of updating the parameters of a value function approximation or policy from data. An RL agent that “learns” a Q-function is not forming beliefs about an unknown distribution; it is running a stochastic approximation algorithm to find a fixed point of an operator. Finally, function approximation in RL refers to representing value functions or policies using parameterized function classes (linear combinations of basis functions, kernel methods, or neural networks), which econometricians will recognize as sieve estimation or nonparametric series estimation: the approximation of an unknown function by projection onto a finite-dimensional basis.
Structural Equivalences
Beyond terminological differences, several formal objects in RL and economics are mathematically identical. The softmax (or Boltzmann) policy used throughout RL is the multinomial logit model of McFadden (1974). The RL softmax policy selects actions according to
$\pi(a\mid s)=\frac{\exp(Q(s,a)/\tau)}{\sum_{a^{\prime}\in\mathcal{A}}\exp(Q(s,a^{\prime})/\tau)},$ (1)
where $\tau>0$ is a temperature parameter. In the discrete choice framework, $Q(s,a)$ plays the role of the deterministic component of utility $v(a\mid x)$, and $\tau$ is the scale parameter of the Type I extreme value (Gumbel) taste shocks $\varepsilon_{a}$. As $\tau\to 0$, the policy converges to the greedy (deterministic) policy, just as the logit choice probability concentrates on the utility-maximizing alternative as the variance of taste shocks vanishes.
The entropy regularization commonly added to RL objectives is the inclusive value (or log-sum-exp) from the discrete choice literature. The soft value function
$V^{\text{soft}}(s)=\tau\log\sum_{a\in\mathcal{A}}\exp(Q(s,a)/\tau)$ (2)
is identical to the McFadden surplus function $W(x)=\tau\log\sum_{a}\exp(v(a|x)/\tau)+C$, where $C$ is Euler’s constant. In the structural estimation literature, this object appears as the Emax function in dynamic discrete choice models following Rust (1987).
The advantage function $A^{\pi}(s,a)=Q^{\pi}(s,a)-V^{\pi}(s)$ measures how much better action $a$ is compared to the average action under $\pi$. In the discrete choice framework, this is precisely the choice-specific value function net of the ex-ante value function, a quantity that appears in the Hotz and Miller (1993) CCP estimator for dynamic discrete choice models.
Notation
Table 1 maps the notation used throughout this survey to the most common econometric equivalents.
Table 1: Notation mapping between reinforcement learning and economics.
RL Term Symbol Economics Equivalent Symbol
State $s\in\mathcal{S}$ State variable, covariate $x_{t}$, $\Omega_{t}$
Action $a\in\mathcal{A}$ Choice, control variable $d_{t}$, $u_{t}$
Reward $r(s,a)$ Per-period utility, payoff $u(x,d)$
Discount factor $\gamma$ Discount factor $\beta$
Policy $\pi(a|s)$ Decision rule, CCP $P(d|x)$
Value function $V^{\pi}(s)$ Ex-ante value function $\bar{V}_{\theta}(x)$
Q-function $Q^{\pi}(s,a)$ Choice-specific value function $v_{\theta}(x,d)$
Return $G_{t}$ Present discounted value $\sum_{k=0}^{\infty}\beta^{k}u_{t+k}$
Transition $P(s^{\prime}|s,a)$ State transition law $f(x_{t+1}|x_{t},d_{t})$
Learning rate $\alpha$ Step size $\alpha_{n}$
TD error $\delta_{t}$ Bellman residual at sample
@@ 3.2 Board Games
The linear parameterization compresses the value function from $10^{20}$ table entries to dozens of weights, the essential response to the curse of dimensionality. Samuel’s architecture, minimax search to depth $H$ with the learned evaluation function scoring leaves, is an early instance of rollout, where tree search simulates forward, truncating at $H$ and substituting $V(s)$ for exact continuation values. [^8] Bertsekas (2021) interprets this as approximate dynamic programming, where offline training (learning $V$) combined with online planning (tree search) implements a Newton-like step for the Bellman equation. The conceptual apparatus of modern game-playing AI (self-play, evaluation learning, tree search, and function approximation) was present in the 1950s.
@@ 3.3 Optimal Control
Blackwell (1965) established the measure-theoretic foundations for discounted dynamic programming with general state and action spaces. He proved that the Bellman operator $T$ defined by $Tu(s)=\sup_{a}[r(s,a)+\gamma\int u(s^{\prime})P(ds^{\prime}|s,a)]$ is a contraction with modulus $\gamma$: $\|Tu-Tv\|_{\infty}\leq\gamma\|u-v\|_{\infty}$. Banach’s fixed-point theorem then guarantees a unique bounded solution $V^{*}$ to the Bellman equation, with $\|V_{k}-V^{*}\|_{\infty}\leq\gamma^{k}\|V_{0}-V^{*}\|_{\infty}$ under value iteration. The central result concerns stationary policies, which use the same decision rule $f:\mathcal{S}\to\mathcal{A}$ at every period regardless of history. Blackwell proved that if the action space is finite, there exists an optimal stationary policy. For countable action spaces, $\epsilon$-optimal stationary policies exist for every $\epsilon>0$. These results justify the focus on memoryless policies, since optimal behavior depends only on the current state, not on the history of past states and actions. The Bellman equation, Howard’s policy iteration, and Blackwell’s existence theorems constitute the planning framework. Given complete knowledge of $P$ and $r$, these methods compute optimal policies exactly. The challenge of learning without such knowledge is the central problem of reinforcement learning. [^10] For comprehensive treatments of dynamic programming in economics, including numerical methods, computational complexity, and the curse of dimensionality, see Rust (2008) and Rust (1996).
@@ 4.1.1 Monte Carlo Estimation
When $P(s^{\prime}|s,a)$ and $r(s,a)$ are unknown, the obvious approach is to use Monte Carlo (MC) to approximate them. These methods estimate value functions from sampled episodes $(s_{0},a_{0},r_{1},s_{1},a_{1},r_{2},\ldots,s_{T})$. The realized return from state $s_{t}$ is
@@ 4.1.2 Sutton (1988)
The target $r_{t+1}+\gamma V(s_{t+1})$ depends on one random reward and one random transition, so its variance is low. The cost is bias. The bootstrap target $V(s_{t+1})$ is the agent’s current estimate, not the true value. This is “bootstrapping”. [^12] See Section 2 for the distinction between RL bootstrapping and Efron’s resampling procedure. As $V$ improves, the bias shrinks; the low variance persists regardless. Sutton demonstrated this tradeoff on a five-state random walk where TD(0) converged faster than Monte Carlo with less data.
The general TD($\lambda$) update interpolates between these extremes through an eligibility trace. [^13] The eligibility trace records which states were recently visited, allowing credit assignment to propagate backward in time. States visited more recently receive stronger updates when the TD error is observed.
@@ 4.1.3 Watkins (1989)
Q-learning achieves this via the update
@@
Q-learning converges to $Q^{*}$ under standard regularity conditions (Section 5.2). The maximization over $a^{\prime}$ makes Q-learning off-policy: [^15] See Section 2 for the on-policy/off-policy distinction. Q-learning learns about the greedy policy $\pi^{*}(s)=\mathop{\it argmax}_{a}Q(s,a)$ while collecting data with an exploratory policy. This exploratory policy needs only to be sufficiently “exploratory” and admits a wide range of policies; including a fully random policy. the update target uses the greedy action at the next state regardless of the action actually taken. Therefore the agent can follow an $\varepsilon$-greedy exploration strategy or even a fully random policy, while learning about the optimal policy directly.
@@ 4.1.4 Williams (1992)
Value-based methods learn an action-value function and derive a policy from it. Williams (1992) derived an alternative, policy gradients, that optimize the policy directly by gradient ascent on expected return
@@
Consider a robot arm that must apply a continuous torque $a\in\mathbb{R}$ to reach a target angle. Q-learning requires computing $\max_{a}Q(s,a)$ at every update, which becomes a nested optimization problem  [^16] In continuous action spaces, $\max_{a}Q(s,a)$ has no closed-form solution in general and must be solved numerically at every Bellman update. Discretizing a $d$-dimensional action space on a grid of $m$ points per dimension costs $O(m^{d})$ evaluations per update step. when the action space is continuous. A policy gradient method sidesteps the issue. Parameterize the policy as a Gaussian $\pi_{\theta}(a|s)=\mathcal{N}(\mu_{\theta}(s),\,\sigma_{\theta}^{2}(s))$, [^17] The Gaussian distribution $\mathcal{N}(\mu,\sigma^{2})$ has density $(2\pi\sigma^{2})^{-1/2}\exp(-(a-\mu)^{2}/2\sigma^{2})$. Here $\mu_{\theta}(s)$ and $\sigma_{\theta}(s)$ are neural network outputs parameterizing the policy mean and standard deviation. sample an action, observe the return, and update $\theta$ by REINFORCE. Virtually all continuous-control results in reinforcement learning descend from the policy gradient framework for this reason.